air contains 1.3% argon by mass. calculate the mass of liquid air needed to obtain 750000tonnes of argon.
1
Expert's answer
2020-09-01T07:36:04-0400

The mass of liquid air needed to obtain 750000 tonnes of argon is 4.28× 10^(13) kg. m (dry air) = 750000 t ×100%/1,3% = 57692307.7 tonnes

p ( dry air)=1.292 kg/m^3

p (liquid air) = 960 kg/m^3

m (liquid air) =57692307.7×1000×960/1.292=4.28× 10^(13) kg